# Phase vs coin vs position disorder as a probe for the resilience and revival of single particle entanglement in cyclic quantum walks

### v1

@@ Abstract
Introduction.– Single-particle entanglement or SPE refers to entanglement between different degrees of freedom (DoF) of a single particle such as a laser-cooled ion trap or a photon [1, 2, 4, 2, 3, 5]. Quantum states exhibiting SPE can encode more extensive quantum information (QI) than their multi-particle analogs exploiting large Hilbert space [6]. These SPE states are more robust to decoherence and advantageous for quantum communication and cryptography [2, 4, 7]. Further, these have applications in large-scale QI processing, quantum state exploration, and studying the interaction between fundamental particles  [2, 8, 9, 10, 11, 12]. However, these are challenging to generate and control. Such states with large or maximal SPE can be generated efficiently via discrete-time quantum walks (DTQWs) [4, 3, 7, 14, 13]. Note that DTQWs have experimental realization using lattice-based quantum systems whose position space matches the discrete lattice sites, such as cold atoms [15, 16], photons [17, 18, 19, 20, 21], trapped ion [22, 23] and NMR system [24].
@@
Furthermore, SPE can be enhanced with moderate or more significant phase or coin disorder strength at specific time steps. Interestingly, disorder can also revive single-particle entanglement. One can tune the SPE value by controlling the strength of various disorders [26], and this work aligns with lab implementations [17, 18, 19, 20, 21] for quantum systems evolving via CQWs.
Below, we present a generalized CQW model (with and without the disorder) for odd- and even-site cyclic graphs that beget a rich class of quantum states and generate SPE between the walker’s position-coin DoF. Then, we discuss the behavior of generated SPE (from clean CQW dynamics) under phase disorder, coin disorder, and position disorder of arbitrary strengths. After that, we analyze the disorder effects on the generated SPE states, compare our results with other relevant approaches, and finally draw conclusions from this comprehensive study. Additional details on generalized disordered CQW dynamics, its properties, and SPE vs disorder for both 4- and 3-cycles are provided in the Supplementary Material (SM) [27].
@@
Figure 1: (a) Clean-CQW; (b) CQW with phase or coin disorder, where amplitudes in coin basis change; (c) CQW with position disorder where the jump length in shift changes randomly. Sites are shown as green dots, and red (for the same amplitudes) and blue (for different amplitudes) curves denote the amplitudes of the particle on 4-cycle.
@@
where $\alpha_{s}(x)$ with $s\in\{0,1\}$ are site-dependent phase angles. Any CQW property will only depend on the relative phase difference $\alpha_{0}(x)-\alpha_{1}(x)$, as we can factor out a global phase $e^{i\alpha_{1}(x)}$ (while estimating a physical observable). Thus, randomizing the phase angle $\alpha_{0}(x)$ only, i.e., choosing randomly from the set $[0,\delta]$ in units of $\pi$ with a uniform distribution [30] and keeping $\alpha_{1}(x)=0$ (constant), introduces a finite random phase difference, i.e., a generalized phase disorder in the CQW setup. The probability density function for a uniform distribution reads,
@@
Here, $a=0,b=\delta$ and the full CQW evolution is then, $U_{k}(t)=\hat{S}_{sp}\cdot[\sum_{x}\ket{x}\bra{x}\otimes\hat{H}]\;$. Clearly, at $\delta=0$, we retrieve the clean-CQW. Thus,
|$\ket{\psi(t)}=\hat{S}_{sp}\cdot[\sum_{x}\ket{x}\bra{x}\otimes\hat{H}]\ket{\psi(t-1)}\;.$ (7)
Figure 2: (a) SPE $\langle E_{av}\rangle$ vs. time step $t$ for different static-phase-disorder strengths ($\delta$); (b) SPE $\langle E_{av}\rangle$ vs. static-phase-disorder strength $\delta$ for different time-steps $t$ for 4-cycle.
In the numerical simulation of the static-phase disorder case, a set of random on-site phase $\alpha_{0}(x)$ values is generated, with its size corresponding to the maximum spread of the walker for up to a specific time-step (say, $t=15$). This site-dependent set remains constant throughout the specified time steps, representing one disorder realization. In our study, any CQW measure $\Omega_{1}(t)$, such as the entanglement entropy ($E_{av}$) or the walker’s position probability ($P_{av}(x=0)$), is first evaluated for a specific disorder realization. In this study, 500 such disorder realizations are computed to estimate the realization-average (quenched average, as one random-phase set is fixed to a particular realization) [25], $\langle\Omega_{1}(t)\rangle$, i.e.,
$\begin{split}\langle\Omega_{1}(t)\rangle=\frac{1}{500}\sum_{j=1}^{500}\Omega_{1}(\ket{\psi(t)}_{j}\bra{\psi(t)}_{j})\;\;\end{split}$ (8)
@@
$\rho_{c}(t=1)=\text{Tr}_{p}(\ket{\psi(1)}\bra{\psi(1)}),\text{ i.e., }\rho_{c}(t=1)=\frac{1}{2}\begin{pmatrix}1&0\\ 0&1\end{pmatrix},$ (11)
which is maximally mixed independent of the choice of initial state parameter $\theta$, random phase angles $\alpha_{0}(0)$ and the number of sites ($k$) in the cyclic graph. The entanglement entropy of this state is,
@@
which proves the state in Eq. (10) has maximal SPE and is a MESPS (a single-particle Bell state). The MESPS at $t=1$ is robust against static phase-disorder of arbitrary strength and for any initial state; also see Fig. 2(b). The recurrence or periodicity of SPE via 4-cycle CQW is lost for phase disorder, and no recurrence occurs for 3-cycle under such disorder, i.e., it remains chaotic even under such disorder, also see SM Fig. 7. From Fig. 2(b), one can observe that SPE remains almost unaffected by small disorder strengths $\delta$; see, for example, SPE values at $t=3,5,7,...$ and the SPE increases under the phase disorder too, see for example SPE values at $t=6,10,14,...$. In addition, we observe that the static phase disorder can revive single-particle entanglement from zero, i.e., from an otherwise separable state observed at $t=4,8,12$, see Fig. 2(b) (cyan, yellow, grey) and Fig. 2(a) (blue). We find similar results for the 3-cycle case, see SM (Fig. 7), i.e., SPE can increase or decrease for static phase disorder of appropriate strength. Moreover, at $t=1$, MESPS is resilient to static phase disorder and initial state choice, as seen in the 4-cycle case.
Dynamic-phase disorder.– The dynamic-phase disorder can be introduced in the CQW setup through the following translation operator (from Eq. (4)), i.e.,
$\begin{split}\hat{S}_{dp}(t)=&\sum_{s=0}^{1}\sum_{x=0}^{k-1}\ket{((x+2s-1)\text{ mod }k)_{p}}\bra{x}\;e^{i\beta_{s}(t)}\otimes\ket{s_{c}}\bra{s_{c}},\end{split}$ (13)
where the time-dependent relative phase angles $\beta_{0}(t)$ (with $\beta_{1}(t)=0$ being global, for the similar reason discussed below Eq. (5)) are random numbers $[0,\delta]$ in units of $\pi$ drawn from uniform distribution, i.e., Eq. (6) with $\beta_{0}(t)$ in the place of $\alpha_{0}(x)$. The full CQW evolution is now, $U_{k}(t)=\hat{S}_{dp}(t)\cdot[\sum_{x}\ket{x}\bra{x}\otimes\hat{H}]\;$ (which is different from $U_{k}(t)$ for static-phase disorder). Finally,
@@
In the numerical simulation of the dynamic disorder, unlike the static-phase case, a set of time-dependent (and site-independent) random phase $\beta_{0}(t)$ angles is generated, with its length equal to the maximum time steps considered for the CQW dynamics (say, $t=15$). It represents one disorder realization, and subsequently, a realization average is computed with 500 such realizations for any quantity $\Omega_{2}(t)$, such as $E_{av}$ or $P_{av}(x=0)$, i.e.,
$\begin{split}\langle\Omega_{2}(t)\rangle=\frac{1}{500}\sum_{j=1}^{500}\Omega_{2}(\ket{\psi(t)}_{j}\bra{\psi(t)}_{j})\;.\end{split}$ (15)
A dynamic phase disorder preserves both the odd-even parity for a 4-cycle and the particle’s physical time-cone for all cycles, as also seen for the static variant; see SM Fig. 6 (red).
Figure 3: (a) SPE $\langle E_{av}\rangle$ vs. time step $t$ for different dynamic-phase-disorder strength ($\delta$); (b) SPE $\langle E_{av}\rangle$ vs. dynamic-phase-disorder strength $\delta$ for different time-steps $t$ for 4-cycle.
@@
Coin disorder and SPE in CQW.– Coin disorder in CQW setup can also be site-dependent (static) or time-dependent (dynamic). See SM Sec. C for the detailed implementation and analysis. For both static and dynamic coin disorder of arbitrary strength, MESPS at $t=1$ remains unperturbed in both 3- and 4-cycles, for two phase-symmetric initial states, i.e., Eq. (1) with $\theta=\frac{\pi}{2}$ with $\phi=\frac{\pi}{2},\frac{3\pi}{2}$, i.e., |$\frac{\ket{0_{p},0_{c}}\pm i\ket{0_{p},1_{c}}}{\sqrt{2}}$. See SM Sec. C for its detailed analytical derivation. Therein, we also observe that SPE remains almost unaffected by low levels of coin disorder. The static and dynamic coin disorders can also increase SPE at specific time steps. In addition, we observe that both the static and dynamic coin disorders revive SPE from zero, i.e., from an otherwise separable state observed at $t=4,8,12$, see SM Sec. C and Fig. 9(b). Static and dynamic coin disorder have many similar effects on CQW and generated SPE values, such as both preserving odd-even parity (4-cycle), obeying physical time-cone (all cycles), increasing SPE (at specific time steps $t$) as well as reviving SPE from zero. However, static coin disorder, on average, is the more substantial coin disorder, see SM Sec. C and Figs. 9-12, where we see the static coin disorder reduces SPE more than the dynamic variant for both 3- and 4-cycles.
@@
$\begin{split}\langle\Omega_{5}(t)\rangle=\frac{1}{500}\sum_{j=1}^{500}\Omega_{5}(\ket{\psi(t)}_{j}\bra{\psi(t)}_{j})\;.\;\end{split}$ (16)
Figure 4: (a) SPE $\langle E_{av}\rangle$ vs. time step $t$ for different position-disorder strength ($\lambda$); (b) SPE $\langle E_{av}\rangle$ vs. position-disorder strength $\lambda$ for different time-steps $t$ for 4-cycle.
@@
Robust MESPS (Bell state) at $t=1$ to disorder
@@
We have also compared our results with prior works, see Refs. [25, 26, 37, 30, 33] on DTQW with 1D line in SM Table 2 and SM Sec. F. Refs. [25, 26, 37, 30, 33] focus on generating SPE using disorder (these works considered either a coin or phase or position disorder) and that too for some particular initial states. These proposals do not look at the effects of a disorder on MESPS or SPE (via clean QW). We bridge this gap and comprehensively report on generating SPE via disorder. We also discuss the effects of various disorders on SPE’s enhancement, revival, and resilience under phase, coin, and position disorder of arbitrary strength with a most general initial state.
Conclusion and outlook.– This study explores the robustness, possibility of revival, and vulnerability of single-particle entanglement (SPE) under phase, coin, and position disorder. Disorders destroy the recurrence of SPE in CQW. The resilience of SPE to minor disorders, particularly phase and coin disorder, demonstrates its potential for stable large-scale QI processing. Phase disorder, especially, leaves maximal SPE (MESPS) states (at $t$=1) completely unperturbed for any disorder strength, irrespective of the choice of the initial state. In contrast, phase-symmetric initial states are entirely immune to coin disorder of arbitrary strength. Moreover, phase and coin disorder enhances SPE at specific time steps. They also revive SPE from zero, i.e., from an otherwise separable state, irrespective of the choice of the initial state.
However, unlike coin or phase disorder, position disorder is more deleterious, breaking the odd-even parity and distorting the walker’s physical time cone. Thus, position disorder generates rich quantum states, which are helpful in quantum cryptography in designing cryptographic keys, as these states will be more challenging to decipher by eavesdroppers as compared to trivial states (product or symmetrically superposed states) [4, 7, 32]. Besides, position disorder can enhance particle dispersion, increase position standard deviation, and be a resource for an efficient quantum search algorithm and analyzing transport phenomena [34, 35, 33, 36]. Interestingly, despite its disruptive nature, position disorder causes SPE to converge to a fixed value at large $t$, indicating a level of robustness of SPE even under significant disturbances. These findings underline the nuanced role disorder plays in the evolution of SPE, offering valuable insights for optimizing quantum communication, dynamics of intricate quantum systems in a lab setting, and cryptographic protocols that leverage single-particle entangled states.
@@
Herein, we set $\beta_{1}(t=1)=0$ for being the global phase factor, as mentioned in the main text on page 4. The evolved state Eq. (21) at $t=1$, is then,
@@
.5 Algorithm to simulate the effects of disorder on single particle entanglement in a cyclic quantum walk

### v2

@@ Abstract
Introduction.– Single-particle entanglement or SPE refers to entanglement between different degrees of freedom (DoF) of a single particle such as a laser-cooled ion trap or a photon [1, 2, 4, 2, 3, 5]. Quantum states exhibiting SPE can encode more extensive quantum information (QI) than their multi-particle analogs exploiting large Hilbert space [6]. These SPE states are more robust to decoherence and advantageous for quantum communication and cryptography [2, 4, 7]. Further, these have applications in large-scale QI processing, quantum state exploration, and studying the interaction between fundamental particles  [2, 8, 9, 10, 11, 12]. However, these are challenging to generate and control. Such states with large or maximal SPE can be generated efficiently via discrete-time quantum walks (DTQWs) [4, 3, 7, 13, 15, 14]. Note that DTQWs have experimental realization using lattice-based quantum systems whose position space matches the discrete lattice sites, such as cold atoms [16, 17], photons [18, 19, 20, 21, 22], trapped ion [23, 24] and NMR system [25].
@@
Further, SPE can be enhanced with moderate or more significant phase or coin disorder strength at specific time steps. Notably, disorder can also revive single-particle entanglement. One can tune the SPE value by controlling the strength of various disorders [27], and this work aligns with lab implementations [18, 19, 20, 21, 22] for quantum systems evolving via CQWs. This counterintuitive result of enhancement, revival and resilience of SPE is a byproduct of the interplay between disorder, interference and unitary quantum evolution [31, 29, 28, 30]. This motivates strategies to harness noise/disorder for enhancing entanglement, safeguarding quantum coherence, and exploring quantum error mitigation (which is reliant on precise noise characterization) techniques for large-scale QI processing and quantum computation [31, 34, 33, 32].
Below, we present a generalized CQW model (with and without the disorder) for odd- and even-site cyclic graphs that beget a rich class of quantum states and generate SPE between the walker’s position-coin DoF. Then, we discuss the behavior of generated SPE (from clean CQW dynamics) under phase disorder, coin disorder, and position disorder of arbitrary strengths. After that, we analyze the disorder effects on the generated SPE states, compare our results with other relevant approaches, and finally draw conclusions from this comprehensive study. Additional details on generalized disordered CQW dynamics, its properties, and SPE vs disorder for both 4- and 3-cycles are provided in Supplementary Material (SM) [35].
@@
Figure 1: (a) Clean-CQW; (b) CQW with phase or coin disorder, where amplitudes in coin basis change; (c) CQW with position disorder where jump length in shift changes randomly. Sites are shown as green dots, and red (for the same amplitudes) and blue (for different amplitudes) curves denote amplitudes of the particle on 4-cycle.
@@
where $\alpha_{s}(x)$ with $s\in\{0,1\}$ are site-dependent phase angles. Any CQW property will only depend on the fluctuations in relative phase difference $\alpha_{0}(x)-\alpha_{1}(x)$, as we can factor out the phase $e^{i\alpha_{1}(x)}$ (while estimating a physical observable), i.e.,
$\hat{S}_{sp}=\sum_{x=0}^{k-1}e^{i\alpha_{1}(x)}(\ket{((x-1)\text{ mod }k)_{p}}\bra{x}\;e^{i(\alpha_{0}(x)-\alpha_{1}(x))}\\ \otimes\ket{0_{c}}\bra{0_{c}}+\ket{((x+1)\text{ mod }k)_{p}}\bra{x}\;\otimes\ket{1_{c}}\bra{1_{c}}).$
Thus, randomizing the phase angle $\alpha_{0}(x)$ only, i.e., by choosing it randomly from the set $[0,\delta]$ in units of $\pi$ with a uniform distribution [38], and keeping $\alpha_{1}(x)=0$ (constant), introduces a finite random phase difference, i.e., a generalized phase disorder in the CQW setup. The probability density function for a uniform distribution reads,
@@
Here, $a=0,b=\delta$ and the full CQW evolution is then, $U_{k}(t)=\hat{S}_{sp}\cdot[\sum_{x}\ket{x}\bra{x}\otimes\hat{H}]\;$. Clearly, at $\delta=0$, we retrieve the clean-CQW. Thus, |$\ket{\psi(t)}=\hat{S}_{sp}\cdot[\sum_{x}\ket{x}\bra{x}\otimes\hat{H}]\ket{\psi(t-1)}\;.$
Figure 2: (a) SPE $\langle E_{av}\rangle$ vs. time step $t$ for different static-phase-disorder strengths ($\delta$); (b) SPE $\langle E_{av}\rangle$ vs. static-phase-disorder strength $\delta$ for different $t$ for 4-cycle.
In the numerical simulation of the static-phase disorder case, a set of random on-site phase $\alpha_{0}(x)$ values is generated, with its size corresponding to the maximum spread of the walker for up to a specific time-step (say, $t=15$). This site-dependent set remains constant throughout the specified time steps, representing one disorder realization. In our study, any CQW measure $\Omega_{1}(t)$, such as the entanglement entropy ($E_{av}$) or the walker’s position probability ($P_{av}(x=0)$), is first evaluated for a specific disorder realization. In this study, 500 such disorder realizations are computed to estimate the realization-average (quenched average, as one random-phase set is fixed to a particular realization) [26], $\langle\Omega_{1}(t)\rangle$,
i.e., $\langle\Omega_{1}(t)\rangle=\frac{1}{500}\sum_{j=1}^{500}\Omega_{1}(\ket{\psi(t)}_{j}\bra{\psi(t)}_{j}).\;\;$
@@
$\rho_{c}(t=1)=\text{Tr}_{p}(\ket{\psi(1)}\bra{\psi(1)}),\text{ i.e., }\rho_{c}(t=1)=\frac{1}{2}\begin{pmatrix}1&0\\ 0&1\end{pmatrix},$ which is maximally mixed independent of the choice of initial state parameter $\theta$, random phase angles $\alpha_{0}(0)$ and the number of sites ($k$) in the cyclic graph. The entanglement entropy of this state is,
@@
which proves the state in Eq. (8) has maximal SPE and is a MESPS (a single-particle Bell state). The MESPS at $t=1$ is robust against static phase-disorder of arbitrary strength and for any initial state, i.e., for an infinite number of initial states (Eq. (1)) with arbitrary $\theta$; also see Fig. 2(b). The recurrence or periodicity of SPE via 4-cycle CQW is lost for phase disorder, and no recurrence occurs for 3-cycle under such disorder, i.e., it remains chaotic even under such disorder, also see SM Fig. 3. From Fig. 2(b), one can observe that SPE remains almost unaffected by small disorder strengths $\delta$; see, for example, SPE values at $t=3,5,7,...$. These results are valid for small time-steps ($t\leq 8$), see Fig. 2(a)-(b); for larger $t$, SPE is vulnerable to all disorder strengths due to the strong coupling between the evolved state and disorder, i.e., the disorder effect piles up in course of time and at large $t$, disorder effects on SPE can be quite significant. Further, SPE increases under the phase disorder too, see for example SPE values at $t=6,10,14,...$, in Fig. 2(a)-(b). In addition, we observe that the static phase disorder can revive single-particle entanglement from zero, i.e., from an otherwise separable state observed at $t=4,8,12$, see Fig. 2(b) (cyan, yellow, grey) and Fig. 2(a) (blue). We find similar results for the 3-cycle case, see SM (Fig. 3), i.e., SPE can increase or decrease for static phase disorder of appropriate strength. Moreover, at $t=1$, MESPS is resilient to static phase disorder and initial state choice, as seen in the 4-cycle case.
Dynamic-phase disorder.– The dynamic-phase disorder can be introduced in the CQW setup through the following translation operator (from Eq. (4)), i.e., $\hat{S}_{dp}(t)=\sum_{s=0}^{1}\sum_{x=0}^{k-1}\ket{((x+2s-1)\text{ mod }k)_{p}}\bra{x}\;e^{i\beta_{s}(t)}\otimes\ket{s_{c}}\bra{s_{c}},$ where the time-dependent relative phase angles $\beta_{0}(t)$ (with keeping $\beta_{1}(t)=0$, for the similar reason discussed below Eq. (5)) are random numbers $[0,\delta]$ in units of $\pi$ drawn from uniform distribution, i.e., Eq. (6) with $\beta_{0}(t)$ in the place of $\alpha_{0}(x)$. The full CQW evolution is now, $U_{k}(t)=\hat{S}_{dp}(t)\cdot[\sum_{x}\ket{x}\bra{x}\otimes\hat{H}]\;$ (which is different from $U_{k}(t)$ for static-phase disorder). Finally,
@@
In the numerical simulation of the dynamic disorder, unlike the static-phase case, a set of time-dependent (and site-independent) random phase $\beta_{0}(t)$ angles is generated, with its length equal to the maximum time steps considered for the CQW dynamics (say, $t=15$). It represents one disorder realization, and subsequently, a realization average is computed with 500 such realizations for any quantity $\Omega_{2}(t)$, such as $E_{av}$ or $P_{av}(x=0)$,
i.e., $\langle\Omega_{2}(t)\rangle=\frac{1}{500}\sum_{j=1}^{500}\Omega_{2}(\ket{\psi(t)}_{j}\bra{\psi(t)}_{j})\;.$ A dynamic phase disorder preserves both the odd-even parity for a 4-cycle and the particle’s physical time-cone for all cycles, as also seen for the static variant; see SM Fig. 2 (red).
Figure 3: (a) SPE $\langle E_{av}\rangle$ vs. time step $t$ for different dynamic-phase-disorder strength ($\delta$); (b) SPE $\langle E_{av}\rangle$ vs. dynamic-phase-disorder strength $\delta$ for different $t$ for 4-cycle.
@@
Coin disorder and SPE in CQW.– Coin disorder in CQW setup can also be site-dependent (static) or time-dependent (dynamic). The former is introduced in the CQW setup through modified evolution, i.e., $U_{k}(t)=\hat{S}(t)\cdot[\sum_{x}\ket{x}\bra{x}\otimes\hat{C}(\rho(x))]\;$ with site-dependent coin operators [26] and the latter via, $U_{k}(t)=\hat{S}(t)\cdot[\mathbf{I}_{P}\otimes\hat{C}(\rho(t))]\;$, with time-dependent coin operators. For both static and dynamic coin disorder of strength ($\omega$), translation operator is $\hat{S}(t)$ (see, Eq. 4), with hopping length $J(t)=1,\forall t$, and MESPS at $t=1$ remains unperturbed for both 3- and 4-cycles, for two phase-symmetric initial states, i.e., |$\frac{\ket{0_{p},0_{c}}\pm i\ket{0_{p},1_{c}}}{\sqrt{2}}$, see Eq. (1) with $\theta=\frac{\pi}{2}$ with $\phi=\frac{\pi}{2},\frac{3\pi}{2}$. See SM Sec. C for its complete analytical derivation and more details. Therein, we also observe that SPE remains almost unaffected by low levels of coin disorder. The static and dynamic coin disorders can also increase SPE at specific time steps. In addition, we observe that both the static and dynamic coin disorders revive SPE from zero, i.e., from an otherwise separable state observed at $t=4,8,12$, see SM Sec. C and Fig. 5(b). Static and dynamic coin disorder have many similar effects on CQW and generated SPE values, such as both preserving odd-even parity (4-cycle), obeying physical time-cone (all cycles), increasing SPE (at specific time steps $t$) as well as reviving SPE from zero. However, static coin disorder, on average, is the more substantial coin disorder, see SM Sec. C and Figs. 5-8, where we see the static coin disorder reduces SPE more than the dynamic variant for both 3- and 4-cycles.
@@
Figure 4: (a) SPE $\langle E_{av}\rangle$ vs. time step $t$ for different position-disorder strength ($\lambda$); (b) SPE $\langle E_{av}\rangle$ vs. position-disorder strength $\lambda$ for different $t$ for 4-cycle.
@@
Robust MESPS (Bell state) at $t=1$ to
disorder
@@
We have also compared our results with prior works, see Refs. [26, 27, 45, 38, 41] on DTQW with 1D line in SM Table I and SM Sec. F. Refs. [26, 27, 45, 38, 41] focus on generating SPE using disorder (these works considered either a coin or phase or position disorder) and that too for some particular initial states. These proposals do not look at the effects of a disorder on MESPS or SPE (via clean QW). We bridge this gap and comprehensively report on generating SPE via disorder. We also discuss the effects of various disorders on SPE’s enhancement, revival, and resilience under phase, coin, and position disorder of arbitrary strength with a most general initial state.
This proposal can be experimentally realized using single-photons with optical elements (e.g., polarizing beam splitters, waveplates, Jones plates), wherein the photon’s polarization encodes the coin state and time bins or orbital angular momentum encodes the position state [18, 49, 48, 47, 46].
Conclusion and outlook.– This study explores the robustness, possibility of revival, and vulnerability of single-particle entanglement (SPE) under phase, coin, and position disorder. Disorders destroy the recurrence of SPE in CQW. The resilience of SPE to minor disorders, particularly phase and coin disorder, demonstrates its potential for stable large-scale QI processing. Phase disorder, especially, leaves maximal SPE (MESPS) states (at $t$=1) completely unperturbed for any disorder strength, irrespective of the choice of the initial state, i.e., for any arbitrary $\theta$ in Eq. (1). Further, phase-symmetric initial states are entirely immune to coin disorder of arbitrary strength. Moreover, phase and coin disorder enhances SPE at specific time steps. These also revive SPE from zero, i.e., from an otherwise separable state, irrespective of the choice of the initial state.
However, unlike coin or phase disorder, position disorder is more deleterious, breaking the odd-even parity and distorting the walker’s physical time cone. Position disorder thus generates rich quantum states, which are helpful in quantum cryptography in designing cryptographic keys, as these states will be more challenging to decipher by eavesdroppers as compared to trivial states (product or symmetrically superposed states) [4, 7, 40]. Besides, position disorder can enhance particle dispersion, increase position standard deviation, and be a resource for an efficient quantum search algorithm and in analyzing transport phenomena [42, 43, 41, 44]. Interestingly, despite its disruptive nature, position disorder causes SPE to converge to a fixed value at large $t$, indicating a level of robustness of SPE even under significant disturbances. These findings underline the nuanced role disorder plays in the evolution of SPE, offering valuable insights for optimizing quantum communication, dynamics of intricate quantum systems in a lab setting, and cryptographic protocols that leverage single-particle entangled states.
@@
Herein, we set $\beta_{1}(t=1)=0$, as mentioned in the main text on page 4. The evolved state Eq. (15) at $t=1$, is then,
@@
.5 Algorithm to see effects of disorders via CQW on SPE